a line having a slope of 3/4 passes through the point(-8,4) . Write the equation of this line in slope-intercpt form.
Answer by mananth(16946)

slope of 3/4 passes through the point(-8,4)
y = mx+b
plug the value of the point & slope to get b
4=-8*3/4 +b
4=-6+b
b=10
the equation = y = 3x/4+10